I need to solve this by completing the square..not the quad. formula
7x squared + 3x -4=0
Thanks for your help!

Answer by sudhanshu_kmr(1152)


7x squared + 3x -4=0
=> 7x2+ 3x -4 = 0
=> 7x2+ 7x - 4x -4 = 0
=> 7(x+1) -4(x+1) = 0
=> (7x-4)(x+1) = 0
x = 4/7 or x = -1